OpenStudy (anonymous):

Find the expression for g(x)? When f(x)=1/(x-5) and g(x) =x^2+2?

OpenStudy (aum):

Find the expression for g(x)? g(x) is given in the problem. Are you sure it is g(x)?

OpenStudy (anonymous):

yeah. I didn't understand the problem either, but that's what is there.

OpenStudy (aum):

Could you post a screenshot of the question?

OpenStudy (anonymous):

it's on a piece of paper :/

OpenStudy (aum):

If you question is correct then g(x) = x^2 + 2. f(x) will not be used in this problem. My guess is it is probably find the expression for f(g(x)) or (fog)(x)

OpenStudy (anonymous):

ok

OpenStudy (aum):

If it is f(g(x)), then To find f(g(x)) replace x on the right hand side of f(x) by g(x) which is x^2 + 2 f(x) = 1 / (x - 5) f(g(x)) = f(x^2+2) = 1 / (x^2+2-5) = 1/(x^2-3)

OpenStudy (anonymous):

thank you

OpenStudy (aum):

you are welcome.